# Split Two-Periodic Aztec Diamond

### v3

@@ 2.1. An Integral Representation of the Correlation Kernel
(16) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m<m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi-\xi^{\prime}}\Phi_{\beta}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
when $\nicefrac{{N}}{{2}}<m^{\prime}<N$. Here $\gamma_{1}$ is a contour surrounding $1$ and $\gamma_{0,1}$ is a contour surrounding 00 and $\gamma_{1}$. Both contours are oriented in the positive direction. In both cases if $m\leq\nicefrac{{N}}{{2}}$ then $\varepsilon=\alpha$ and if $m>\nicefrac{{N}}{{2}}$ then $\varepsilon=\beta$.
@@
Figure 6 depicts the contours from the above formula. The proof of Theorem 2.1 is in Section 3, with some details deferred to Sections 5 and 6.
The kernel presented in Theorem 2.1 has somewhat different forms depending on which side of the interface the vertex $(4m^{\prime},2\xi^{\prime}+j)$ is on, however this is simply due to the way the kernel is derived. The following corollary illustrates that there is a symmetry to the correlation kernel when we flip the model about the interface.
Corollary 2.1.
For $\nicefrac{{N}}{{2}}<m^{\prime}<N$ the correlation kernel of the split two periodic Aztec diamond may be written as,
(17) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\beta}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,2}(w)F_{\alpha,1}(w)F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,1}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(z)F_{\alpha,2}(z)F_{\beta,2}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
We us the notation $\gamma_{\text{Br}}$ to denote a contour surrounding the branch cuts $(-\infty,-\delta^{-2}]\cup[-\delta^{2},0]$ in the clockwise direction, where $\delta=\max\{\alpha,\beta\}$
This corollary is proven at the end of Section 3. Despite the fact that the initial form of the kernel was more compact, this expression of the kernel is more useful for making comparisons to the typical two-periodic model, which we consider below in Section 2.2. As was done in Corollary 2.1, we can modify the second double contour integral in (15) so that the contour $\gamma_{1}$ is deformed to the contour $\gamma_{\text{Br}}$. In doing so, we also get a single contour integral term of a slightly different form,
$\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ =\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,2}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
For brevity, we define
(18) $\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\\ \frac{\mathbb{I}_{m^{\prime}\leq\frac{N}{2}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,2}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{\mathbb{I}_{m^{\prime}>\frac{N}{2}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,1}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(z)F_{\alpha,2}(z)F_{\beta,2}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
Now we may succinctly state the kernel for any $0<m^{\prime}<N$,
(19) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\varepsilon(m^{\prime})}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon(m)}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon(m^{\prime}),1}(w)^{\frac{N}{2}-m^{\prime}}F_{\varepsilon(m^{\prime}),1}(w)\Phi_{\varepsilon(m)}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon(m^{\prime}),2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2}{1+2g_{\alpha,\beta}(w)}\\ \times F_{\varepsilon(m^{\prime}),2}(w)F_{\overline{\varepsilon}(m^{\prime}),1}(w)F_{\varepsilon(m^{\prime}),1}(w)\Phi_{\varepsilon(m)}(z)^{m-\frac{N}{2}}\\ +\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}$
Here we use the following notation,
$\big{(}\varepsilon(m),\overline{\varepsilon}(m)\big{)}=\begin{cases}(\alpha,\beta)&\text{if }m\leq\nicefrac{{N}}{{2}}\\ (\beta,\alpha)&\text{if }m>\nicefrac{{N}}{{2}}\end{cases}$
In instances where we know that both vertices are on the same side of the interface, we use the shorthand notation,
$\big{(}\varepsilon,\overline{\varepsilon}\big{)}=\begin{cases}(\alpha,\beta)&\text{if }m^{\prime},m\leq\nicefrac{{N}}{{2}}\\ (\beta,\alpha)&\text{if }m^{\prime},m>\nicefrac{{N}}{{2}}\end{cases}$
@@ 2.2. Comparison to the Two-Periodic Aztec Diamond
Before we analyze the behavior of the model, we first want to comment on the form of the correlation kernel in (19) and how it relates to the correlation kernel of the typical two-periodic Aztec diamond. We recover the typical two-periodic Aztec diamond from the split two-periodic Aztec diamond by setting $\alpha=\beta=\varepsilon$. The correlation kernel of the two-periodic Aztec diamond, as formulated in [3], can be stated using the notation presented in this paper as
@@
(21) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\\ \Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}+\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon,2}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon,k}(z)^{m-\frac{N}{2}}\\ \times\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,2}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,1}(w)F_{\varepsilon,k}(z)\$
From this expression we see how the correlation kernel of the split model relates to the typically two-periodic model. In order to understand how the local asymptotics of the split two-periodic Aztec diamond differ from the local asymptotics of the typical two-periodic Aztec diamond we only need to understand the asymptotics of the additional two double contour integrals and the single integral term $J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)$. These will be our main focus when describing the local asymptotics in Section 2.4.
@@ 2.3. Classification of the Macroscopic Regions
This is the same definition of regions presented for the typical two-periodic model in [11, Definition 2.9]. The saddle function, $\Psi(z)$, can also be used to determine the boundaries between the macroscopic regions. Any point where saddle points coalesce and there is a double or triple saddle point is on some boundary between the macroscopic regions. Additionally, due to the piece wise nature of the saddle functions, we must check for transitions along the line $x=\nicefrac{{1}}{{2}}$. Figure 7 depicts an example of the macroscopic boundaries for a split two-periodic model.
@@ 2.4. Asymptotic Behavior
In this work, we will focus our asymptotic results to the local behavior of the model away from the interface. We define the coordinates in the following manner,
@@
where we assume $x\neq\nicefrac{{1}}{{2}}$, and $x_{1}$ and $x_{2}$ are small enough in magnitude so that the coordinates $(4m,2\xi+i)$ and $(4m^{\prime},2\xi^{\prime}+j)$ lie on the same side of the interface. Under these considerations, we rewrite correlation kernel in terms of the saddle functions,
(26) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\\ \Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}+\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,2}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,1}(w)F_{\varepsilon,k}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,2}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
The saddle function $\psi_{k}(z)$ is defined in equation (23) and we define the function $\varphi_{\varepsilon,k}(z;x_{1},y_{1})$ as
@@
(28) $\Big{[}I^{\varepsilon}_{2,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,2}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,1}(w)\\ \times F_{\varepsilon,k}(z)\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,2}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
@@
(29) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ \Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}+\sum_{k=1,2}\Big{[}I^{\varepsilon}_{2,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}$
Thus the local asymptotics of the split model depends on the asymptotics of the typically two-periodic Aztec diamond along with the asymptotics of the contour integrals $J$, $I^{\varepsilon}_{2,1}$ and $I^{\varepsilon}_{2,2}$. We will start with a statement about $J$.
Proposition 2.1.
For $(m^{\prime},\xi^{\prime})$ and $(m,\xi)$ defined by equations (24) and (25), we have
(30) $|J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq c_{2}e^{-c_{1}N}$
where $c_{1}$ and $c_{2}$ are positive constants.
We will utilize the symmetry of the double contour integrals and just consider the asymptotics on the $\alpha$-side of the interface, e.g. when $x<\nicefrac{{1}}{{2}}$. In doing so, we should consider two cases:
@@
The behavior of the double contour integrals $I_{2,2}^{\alpha}$ does not depend on the macroscopic region of the model, however, if we are in case (b), the decay of $I_{2,2}^{\alpha}$ does depend on the region of the model in other ways. Consider the regions of the model where $(x,y)$ satisfies one of the following inequalities:
@@
Since $\psi^{\prime}_{\alpha,2}(z)$ is linear in $x$ and $y$, the inequalities just denote trapezoidal regions of the model. Note that these regions are empty if $\beta<\alpha$. The exact size and shape of these regions depends on the exact values of $\alpha$ and $\beta$. Figure 8 gives examples of these regions for different values of $\alpha$ and $\beta$.
Figure 8. Regions where $I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)$ decays non-exponentially are shaded in red. On the left side figure, $\beta=\nicefrac{{1}}{{2}}$ and $\alpha=\nicefrac{{1}}{{3}}$, while on the right side figure, $\beta=\nicefrac{{2}}{{3}}$ and $\alpha=\nicefrac{{1}}{{5}}$.
We will denote the asymptotic region where either (i) or (ii) is satisfied as the strong-coupling region of the model, denoted $\mathfrak{C}$. We state the following asymptotic results regarding the behavior of $I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)$,
Proposition 2.2.
$|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq c_{2}e^{-c_{1}N}\;\text{ or }\;|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq c_{3}N^{-1}$
where $c_{1}$, $c_{2}$, and $c_{3}$ are positive constants. The latter occurring when only when $(x,y)\in\mathfrak{C}$.
The behavior of the second double contour integral, $I_{2,1}^{\alpha}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)$, depends on the macroscopic region of the model and also on the strong-coupling region described above. We focus our attention on the behavior of this integral in the rough and smooth regions of the model only. We state the following results,
@@
    $|I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq cN^{-1}$
@@
  3. (3)
    If $(x,y)\in\mathfrak{S}$ and $(x,y)\in\mathfrak{C}$, then
    $|I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq cN^{-1}$
    for some positive constant $c$.

### v4

@@ 2.1. An Integral Representation of the Correlation Kernel
(16) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\beta}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(w)F_{\alpha,2}(w)F_{\beta,2}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
when $\nicefrac{{N}}{{2}}<m^{\prime}<N$. For both cases, if $m\leq\nicefrac{{N}}{{2}}$ then $\varepsilon=\alpha$ and if $m>\nicefrac{{N}}{{2}}$ then $\varepsilon=\beta$. Additionally, $\gamma_{1}$ is a contour surrounding $1$ and $\gamma_{0,1}$ is a contour surrounding 00 and $\gamma_{1}$. Both contours are oriented in the positive direction.
@@
The proof of Theorem 2.1 is in Section 3, with some details deferred to Sections 5 and 6. The kernel presented in Theorem 2.1 is written in a way that makes comparisons to the typical two-periodic Aztec diamond more immediate. There is a way to represent the kernel with only one double contour integral. This form of the kernel is shown in equation (57), which is located in the proof of Theorem 2.1 when $m^{\prime}>\nicefrac{{N}}{{2}}$.
@@ 2.2. Comparison to the Two-Periodic Aztec Diamond
Before we analyze the behavior of the model, we first want to comment on how the form of the correlation kernel relates to the correlation kernel of the typical two-periodic Aztec diamond. From the definition of the model, we can recover the two-periodic Aztec diamond (with two-periodic parameter $\varepsilon$) from the split two-periodic Aztec diamond by setting $\alpha=\beta=\varepsilon$. The correlation kernel of the two-periodic Aztec diamond, as formulated in [3], can be stated using the notation presented in this paper as
@@
At the level of the correlation kernel, we recover the kernel of the typical two-periodic Aztec diamond when we set $\alpha=\beta=\varepsilon$. This is because the second double contour integral presented in equations (15) and (16) will contain the matrix products $F_{\alpha,1}(w)F_{\alpha,2}(w)$ and $F_{\alpha,2}(w)F_{\alpha,1}(w)$, respectively. Both of these products are exactly zero.
@@
(18) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon,l}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon,k}(z)^{m-\frac{N}{2}}\\ \times\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,l}(w)F_{\bar{\varepsilon},\bar{l}}(w)F_{\varepsilon,\bar{l}}(w)F_{\varepsilon,k}(z)\$
where
$(\varepsilon,\bar{\varepsilon})=\begin{cases}(\alpha,\beta)&\text{if }m,\,m^{\prime}<\nicefrac{{N}}{{2}}\\ (\beta,\alpha)&\text{if }m,\,m^{\prime}>\nicefrac{{N}}{{2}}\end{cases}$
and
$(l,\bar{l})=\begin{cases}(2,1)&\text{if }m,\,m^{\prime}<\nicefrac{{N}}{{2}}\\ (1,2)&\text{if }m,\,m^{\prime}>\nicefrac{{N}}{{2}}\end{cases}$
From this expression we see how the correlation kernel of the split model relates to the typically two-periodic model. In order to understand how the local asymptotics of the split two-periodic Aztec diamond differ from the local asymptotics of the typical two-periodic Aztec diamond we only need to understand the asymptotics of the additional two double contour integrals. These will be our main focus when describing the local asymptotics in Section 2.4.
@@ 2.3. Classification of the Macroscopic Regions
This is the same definition of regions presented for the typical two-periodic model in [11, Definition 2.9]. The saddle function, $\Psi(z)$, can also be used to determine the boundaries between the macroscopic regions. Any point $(x,y)$ where saddle points coalesce is on some boundary between the macroscopic regions. Additionally, due to the piece wise nature of the saddle functions, we must check for transitions along the line $x=\nicefrac{{1}}{{2}}$. Figure 7 depicts an example of the macroscopic boundaries for a split two-periodic model.
@@ 2.4. Asymptotic Behavior
Our goal in this section is to summarize the local asymptotics of the model, so we will assume the two coordinates $(4m^{\prime},2\xi^{\prime}+j)$ and $(4m,2\xi+i)$ are on the same side of the interface and that they are asymptotically close. With these considerations in mind, we can start from the phrasing of the kernel given in equation (18).
This work will only discuss the asymptotic results of the model away from the interface. To begin, we define the coordinates in the following manner
@@
where we assume $x\neq\nicefrac{{1}}{{2}}$, and $x_{1}$ and $x_{2}$ are small enough in magnitude so that the coordinates $(4m,2\xi+i)$ and $(4m^{\prime},2\xi^{\prime}+j)$ lie on the same side of the interface. In general, we will assume that $x_{1},\,x_{2},\,y_{1},$ and $y_{2}$ are $o(N)$. Under these considerations, we rewrite correlation kernel in terms of the saddle functions,
(23) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,l}(w)F_{\bar{\varepsilon},\bar{l}}(w)F_{\varepsilon,\bar{l}}(w)F_{\varepsilon,k}(z)\\ \times\exp\left[N\left(\psi_{l}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,l}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
We are using the same definitions of $\varepsilon,\,\bar{\varepsilon},\,l,$ and $\bar{l}$ as defined below equation (18). The saddle functions $\psi_{k}(z)$ are defined by equation (20) and we define the functions $\varphi_{\varepsilon,k}(z;x_{1},y_{1})$ as
@@
(25) $\Big{[}I^{\varepsilon}_{l,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ =\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,l}(w)F_{\bar{\varepsilon},\bar{l}}(w)F_{\varepsilon,\bar{l}}(w)F_{\varepsilon,k}(z)\\ \times\exp\left[N\left(\psi_{l}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,l}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
@@
(26) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\sum_{k=1,2}\Big{[}I^{\varepsilon}_{l,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}$
This work will focus on the asymptotic behavior on the $\alpha$-side of the interface, in other words when $m,\,m^{\prime}<\nicefrac{{N}}{{2}}$. In this case the above equation becomes,
(27) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\alpha}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\sum_{k=1,2}\Big{[}I^{\alpha}_{2,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}$
Ultimately, our goal is to show that the contributions from the contour integrals $I^{\alpha}_{2,1}$ and $I^{\alpha}_{2,2}$ tend to 00 as $N\to\infty$. This will justify the definition of the macroscopic boundary given in Definition 2.1. The order of decay depends on the region of the model of the asymptotic coordinate. We will find that, in certain regions, these additional terms affect the second order term of the local asymptotics of $\mathbb{K}_{N}$.
Before we state the asymptotic results, we will give a brief justification on why we limit our analysis to the $\alpha$-side of the interface. The reason this is justified is due to a symmetry between the saddle functions $\psi_{\varepsilon,1}(z)$ and $\psi_{\varepsilon,2}(z)$. In particular, since $r_{\varepsilon,1}(z)=r_{\varepsilon,2}(z)^{-1}$, this implies
$\psi_{\varepsilon,1}(z;x,y)=\psi_{\varepsilon,2}(z;1-x,y)$
Assuming the coordinate definition given in equations (21) and (22), this means that the local asymptotic behavior of the contour integral $I^{\varepsilon}_{2,2}$ when $x<\nicefrac{{1}}{{2}}$ is the same as the local asymptotic behavior of the contour integral $I^{\varepsilon}_{1,1}$ when $x>\nicefrac{{1}}{{2}}$. [^5] This means if one is $O(N^{-1})$, then so is the other. Constant factors may be differing. The same is true for $I^{\varepsilon}_{2,1}$ and $I^{\varepsilon}_{1,2}$. Instead, to examine all possible cases of model behavior we will consider the following two cases.
@@
This is because The behavior of $I_{2,2}^{\alpha}$ not only depends on the macroscopic regions of the model, but, if we are in case (b) above, it also depends on the region of the model in other ways. Consider the regions of the model where $(x,y)$ satisfies one of the following inequalities:
@@
Since $\psi^{\prime}_{\alpha,2}(z)$ is linear in $x$ and $y$, the inequalities just denote trapezoidal regions of the model. These regions are empty if $\beta<\alpha$. The exact size and shape of these regions depends on the exact values of $\alpha$ and $\beta$. Figure 8 gives examples of these regions for different values of $\alpha$ and $\beta$.
Figure 8. Regions of the model where (i) and (ii) are satisfied are shown in red. We refer to these as the strong-coupling region.
We will denote the asymptotic region where either (i) or (ii) is satisfied as the strong-coupling region of the model, and denote it by $\mathfrak{C}$. We state the following asymptotic results regarding the behavior of $I^{\alpha}_{2,2}$,
@@
    $|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq cN^{-1}$
@@
    $|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq ce^{-c_{1}N}$
    for some positive constants $c$ and $c_{1}$.
  3. (3)
    If $(x,y)\in\mathfrak{S}\cap\mathfrak{C}$, then
    $|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq cN^{-1}$
    for some positive constant $c$.
The behavior of the second double contour integral, $I^{\alpha}_{2,1}$, is less complicated as it does not depend on the macroscopic region of the model. We state the following proposition describing its behavior,
Proposition 2.2.